OpenStudy (secret-ninja):

PLEASE HELP I'LL GIVE A MEDAL The table below represents a linear function f(x) and the equation represents a function g(x): x f(x) -1 -11 0 -1 1 9 g(x) g(x) = 5x + 1 Part A: Write a sentence to compare the slope of the two functions and show the steps you used to determine the slope of f(x) and g(x). Part B: Which function has a greater y-intercept? Justify your answer.

OpenStudy (secret-ninja):

I have both of the slopes for part A, but I don't know how to "compare them", and I still need help for part B.

OpenStudy (igreen):

Easy, the y-intercept for f(x) is -1, because the table states that it has the point (0, -1). The y-intercept is what 'y' is when x is 0. And since the table f(x) says that when x is 0, y is -1, then the y-intercept for f(x) is -1.

OpenStudy (igreen):

Since the y-intercept is what y is when x is 0, let's plug in 0 for x in the equation for g(x): g(x) = 5x + 1 y = 5x + 1 Input 0 for x: y = 5(0) + 1 Multiply: y = 0 + 1 Add: y = 1

OpenStudy (igreen):

f(x)'s y-intercept = -1 g(x)'s y-intercept = 1 1 > -1 Therefore, g(x)'s y-intercept is greater than f(x)'s y-intercept.

OpenStudy (secret-ninja):

So for part B i would put: Part B: The g(x) has a greater intercept because it is not a negative number. right? @iGreen?

OpenStudy (secret-ninja):

Thank you so much!!!! :))))